### v1

# Multi-Objective Neural Network Assisted Design Optimization of Soft Fin-Ray Grippers for Enhanced Grasping Performance

Abstract
Soft Fin-Ray grippers can perform delicate and careful manipulation, which has caused notable attention in different fields. These grippers can handle objects of various forms and sizes safely. The internal structure of the Fin-Ray finger plays a significant role in its adaptability and grasping performance. However, modeling the non-linear grasp force and deformation behaviors for design purposes is challenging. Moreover, when the Fin-Ray finger becomes more rigid and capable of exerting higher forces, it becomes less delicate in handling objects. The contrast between these two objectives gives rise to a multi-objective optimization problem. In this study, we employ finite element method (FEM) to estimate the deflections and contact forces of the Fin-Ray, grasping cylindrical objects. This dataset is then used to construct a multilayer perception (MLP) for prediction of the contact force and the tip displacement. The FEM dataset consists of three input and four target features. The three input features of the MLP and optimization design variables are the thickness of the front and supporting beams, the thickness of the cross beams, and the equal spacing between the cross beams. In addition, the target features are the maximum contact forces and maximum tip displacements in x- and y-directions. The magnitude of maximum contact force and magnitude of maximum tip displacement are the two objectives, showing the trade-off between force and delicate manipulation in soft Fin-Ray grippers. Furthermore, the optimized set of solutions are found using multi-objective optimal techniques. We use non-dominated sorting genetic algorithm (NSGA-II) method for this purpose. Our findings demonstrate that our methodologies can be used to improve the design and gripping performance of soft robotic grippers, helping us to choose a design not only for delicate grasping but also for high-force applications.
Index Terms:
Soft Robotics, Optimization, Neural Networks, Finite Element Method, Fin-Ray Grippers
I Introduction
Robotic grippers used to be made of rigid parts and links in most cases, but recent progress in soft robotics and related fields such as material science has led to attract much attention to soft grippers, and boosting their progress. Soft grippers have the ability to handle a wider range of objects compared to their rigid counterparts, while also allowing for the use of simpler control frameworks[1]. The basic idea of the Fin-Ray grippers has inspired biologically when Leif Kniese observed a phenomenon in fish; by applying force to the structure of the fish’s fins, it bends in the opposite direction of the force. By following up on this observation and studying the structure in collaboration with Rudolf Bannasch, the Fin-Ray Effect was introduced[2]. Fin-Ray grippers can be used to grasp different delicate and sensitive objects such as eggs and fruits[3]. In a study, An et al. designed an optimized gripper based on the Fin-Ray effect with an integrated linkage mechanism to grasp and harvest tomatoes. Their design provided a balanced force distribution on tomatoes and also included a blade at the tip of Fin-Ray finger that automatically cuts the stems while harvesting[4]. FEM is widely considered an accurate method, capable of simulating and modeling linear and multiphysics problems. In the context of soft robotics, it provides the ability to simulate complex structures and nonlinear elastic materials, two essential aspects of understanding soft robotics. However, when dealing with these problems, FEM is computationally expensive and may face convergence difficulties. Combining data-driven methods such as Machine Learning (ML) techniques with FEM offers new opportunities in the modeling and development of soft robotic systems. This combination creates a powerful tool for optimizing soft robotics structures, materials, and control[5]. Crooks et al. designed a novel gripper inspired by the traditional Fin-Ray grippers. Their gripper was activated via a motor-tendon mechanism and made of soft and hard materials, all printable as a complete gripper. Hard parts included crossbeams, supports, and fingernails. They compared their optimized design with the common Fin-Ray grippers both experimentally and by simulation. As a result of their design, gripper’s tip and structure could move more and perform better gripping with a lesser amount of force and actuation compared to the common Fin-Ray grippers. Also, the contact area was greater, causing a more stable grip, and the gripper was able to handle a heavier load[6]. Hazrat Ali et al. did a study on material selection for Fin-Ray grippers. They simulated a problem of single gripper finger’s response to a static force in ANSYS Workbench. They wanted to choose the best material according to their design goals[7]. Shan et al. proposed a kinetostatic model for multi-crossbeam soft fingers to accurately estimate contact forces during object grasping. Their approach not only enabled efficient force prediction but also incorporated the influence of varying stiffness across different finger segments, providing valuable insights for optimizing the design and performance[8]. Elgeneidy et al. used FEM to optimize the structural design of adaptive soft Fin-Ray fingers with the layer jamming mechanism for variable stiffening. This mechanism enabled a low contact force at first, and a higher amount of force in the jammed position because of the angle of ribs. The chosen optimization parameters were rib thickness, the first rib’s angle and the angle increment between ribs. They achieved enhanced shape adaptation and force generation, reduced initial contact forces which is desirable for delicate grasping, and increases final contact forces[9]. Xu et al. introduced a compliant adaptive Fin-Ray gripper. They changed the traditional Fin-Ray gripper by putting rigid parts into it, making its force-deformation behavior linear. Then, they trained a neural network using data from FEM. This made it possible to calculate external force from the finger’s deformation. However, the accuracy of the estimated force dropped as the applied force point moved from the middle to the finger’s ends[10]. De Barrie et al. developed a neural network for real-time prediction of contact forces and stress maps in soft Fin-Ray grippers. They prepared a dataset using FEM simulations of the gripper and different grasped objects, varying in size, shape, height, and angle of approach. Their network demonstrated promising results and was able to predict cases that were not in the training data. However, there were limitations in the real-world applications, including that the network’s performance dropped as the camera angle increased[11]. Deng et al. attempted to find an optimal Fin-Ray finger to improve its grasping. They defined different objects by a superquadratic model and, by considering the winding number, they presented the grasping quality function. They used numerical simulations, and considered thickness and spacing distance as the optimization parameters. They figured out that in real-world applications, computing the grasp quality function is hard due to the lack of precise determination of deformations. Also, the winding number is a geometric property and does not represent any aspect of the control framework[12]. Yao et al. proposed a two dimensional kinetostatic model of a soft Fin-Ray finger. The model was able to calculate the total contact force and deformations, and it was found out that the results obtained within their model exhibited similar outputs and accuracy compared to those of the FEM simulation. However, friction was not considered in the model. Next, they used this model to optimize the finger structure[13]. Yao et al. investigated the internal crossbeam structures of Fin-Ray grippers. They considered four different layouts, including one without crossbeams and used FEM to enhance the adaptability of the Fin-Ray fingers. They found out that removing the internal structure enhanced the finger’s ability to conform to delicate items while minimizing applied force[14]. Although Fin-Ray fingers can take the shape of an object very well, they have low performance in holding heavy loads. Topology optimization is considered a way to improve the grasping ability and safety factor of Fin-Ray fingers. Lakshmi Srinivas et al. compared fingers with different internal layouts both with and without topology optimization. Comparing the best layouts of two groups, while the non-optimized finger offered the best wrapping ability, the optimal one provided a superior balance of wrapping ability, structural strength, and lightweight design, making it an efficient choice for high-load and versatile grasping tasks[15]. Ghanizadeh et al. estimated the contact forces in soft Fin-Ray grippers for open-loop and closed-loop control purposes using FEM and experimental validation. However, this study considers only one internal structure for the Fin-Ray finger which limits the generalization of the findings[16]. Wang et al. proposed a physics-informed neural network (PINN) to model a soft Fin-Ray structure, where the minimum potential energy was integrated into the loss function based on elasticity theory. They trained two models, with and without the data from a real Fin-Ray finger. Their experiments showed that the PINN without the real data and the FEM had nearly same accuracy, while the PINN with the real data stood out and improved the accuracy[17].
This paper presents a method for optimizing the internal structure of the Fin-Ray finger. In the second section, we use FEM to estimate the contact force and tip displacement for the cylinders. In the third section, we employ neural networks to predict our desired target features. In the fourth section, we use a multi-objective optimization technique to find the set of all Pareto efficient solutions. In the last section, we prepare a conclusive summary of the paper.
II Finite Element Method
In this section, we examine the Fin-Ray finger contact forces and tip displacement during interaction with a 20 mm diameter cylinder. We use the data from this part to train the MLP model in the following phases of our study.
We use FEM to estimate the Fin-Ray finger contact force and tip displacement for different internal structures. Static structural analysis is employed to analyze contact forces with the ANSYS software. The cylindrical object with a diameter of 20 mm contacts the Fin-Ray finger at the midpoint of the fronting beam. This cylinder is considered to be linearly elastic with the properties of the acrylonitrile butadiene styrene (ABS) material found in manufacturing technical specifications. Thermoplastic polyurethane (TPU) is widely employed in soft grippers due to its flexibility and resilience[18]. In this study, all simulated Fin-Ray fingers have the same material of TPU 95A, which the standard library of ANSYS materials does not include[11]. Hence, we consider Young’s modulus of 26 MPa and Poisson’s ratio of 0.48 as the mechanical properties of this material[14]. Afterwards, we define the grasp frictional contact region and additional essential constraints for the simulations, such as the fixed displacement constraints. We also use the default mechanically programmed controlled mesh method with an adaptive size function. It is assumed that the Fin-Ray finger experiences a 2.5 cm base displacement in the direction of the cylindrical object. In addition, we conduct a convergence study to make sure that our numerical results are independent of the mesh.
We perform FEM simulations for Fin-ray fingers with different internal structures to determine the tip displacement and estimate the contact force for each case. To this end, we change three design parameters of the Fin-Ray finger internal structure: thickness of the front and supporting beams, the thickness of the crossing beams, and the equal spacing between each crossbeam. The thicknesses of the support and front beams are considered equal in this study. These parameters are vital because of their importance in influencing our target features. The range in which these parameters have changed is summarized in Table I. We simulate all possible structures within this feasible range of thicknesses. Other internal structural parameters, such as the angle of inclination of the crossbeams and the length of the finger base, are maintained unchanged. The front beam and crossbeams are parallel to the y and x axes just before contact, respectively. An illustration of the Fin-Ray finger before and after contact with the cylindrical object is shown in Fig. 1, where $D_{x}$ and $D_{y}$ are the tip displacement of the Fin-Ray finger in the x- and y-directions. Furthermore, the base length, the fronting and supporting beam lengths, and width of the Fin-Ray finger are assumed to be 35 mm, 100 mm, and 30 mm, respectively.
Figure 1: An illustration of the Fin-Ray finger in simulation
TABLE I: Values for each design parameter of Fin-Ray finger
Parameter Name Range Increment Size
Thickness of front and supporting beams 1.5-4 mm 0.5 mm
Thickness of crossbeams 0.6-1.6 mm 0.2 mm
Spacing between each crossbeam 10-20 mm 2 mm
The total displacement contour for one of the internal structures is depicted in Fig. 2 as an example of the results. Parameters for this specific internal structure are the front and supporting beam with thickness of 2 mm, the crossing beams with the thickness of 0.6 mm, and each crossbeam equal spacing of 16 mm.
Figure 2: Total deformation contour for the specified Fin-ray finger in meters
After each FEM simulation, we record values of the maximum contact force and the maximum tip displacement in both x- and y- directions to form a dataset. This dataset is used to train the MLP model in the next phase of the investigation. We note that maximum values of contact force and maximum tip displacement are always achieved at the maximum base displacement of the Fin-Ray finger, which is 2.5 cm. The variation of the contact force with the base displacement is displayed in Fig. 3 for the internal structure with fronting and supporting beam thickness of 2 mm, the crossing beams thickness of 0.6 mm, and each crossbeam equal spacing of 16 mm. For this proposed example, we can see that the force-displacement behavior is almost linear for the first 1.5 cm of base displacement, and in the last 1 cm of base displacement this behavior is nonlinear as a result of the layer jamming effect [19].
Figure 3: Contact force vs. base displacement for the specified Fin-ray finger
III MLP Model
In this section, we derive the MLP model to approximate the behavior of our complex system. This model will be employed later to determine the optimal internal structure of the Fin-Ray finger.
The Min-Max feature scaling is chosen here as the normalization method. We use Pearson correlation heatmap as a graphical tool that displays the correlation between our variables, derived by Eq. 1 where $r$ is the correlation coefficient, $x_{i}$ is the values of the variable $x$ in the sample, and $\bar{x_{i}}$ is the mean of $x$. Also, $y_{i}$ denotes values of the variable $y$ in the sample, and $\bar{y_{i}}$ is the mean of $y$[20]. This heatmap approach suggests that the correlations between each input and target feature are significant, which shows the suitability of the selected features. Other exploratory data analysis (EDA) techniques are used to reduce the effect of any outlier and address similar problems. Hence, we organize the data to ensure accuracy, consistency, and suitability for the task in a proper manner.
${r}={\frac{\sum\left(x_{i}-\bar{x}\right)\left(y_{i}-\bar{y}\right)}{\sqrt{\sum\left(x_{i}-\bar{x}\right)^{2}\sum\left(y_{i}-\bar{y}\right)^{2}}}}$ (1)
A multilayer perceptron (MLP) is chosen to handle complex nonlinear behavior between the input and output features using intermediate hidden layers. In this feedforward neural network, the mean squared error is considered as the loss for the fitted model. The input layer of the MLP has three neurons consisting of three features: the thickness of the front and supporting beams, the thickness of the crossing beams, and the equal spacing between each crossbeam. The output layer consists of four layers, including maximum contact forces and maximum displacements of the tip in the x- and y-directions. The MLP has three hidden and dropout layers. We use 80% of the data for training, 10% for the validation and the remaining 10% for the test. The number of neurons in each hidden layer, the type of activation function for hidden layers, the regularization rate, and the dropout rate are hyperparameters that are searched through the specified range as summarized in Table II. In this search, a specific parameter value has higher scores when it performs better in the validation set. Moreover, the output layer’s activation function is set as Sigmoid function. The final MLP architecture and the selected hyperparameters are summarized in Table III. Subsequently, K-Fold Cross-Validation is used to ensure that our model is trained and tested on representative samples, reducing bias and enhancing the overall performance.
TABLE II: Grid search over specified parameter values
Parameter Name Values
Number of neurons in the first hidden layer 1-2-3-4-5-6-7-8-9-10
Number of neurons in the second hidden layer 1-2-3-4-5-6-7-8-9-10
Number of neurons in the third hidden layer 1-2-3-4-5-6-7-8-9-10
Hidden layers activation function types Leaky ReLU-Sigmoid-Tanh
Regularization rates 0-0.001-0.01
Dropout rates 0-0.1-0.2
TABLE III: Final MLP architecture
Parameter Name Highest Score
Number of neurons in the first hidden layer 9
Number of neurons in the second hidden layer 10
Number of neurons in the third hidden layer 7
Hidden layers activation function type Tanh
Regularization rate 0
Dropout rate 0
The loss curves for the first, last, and mean of all folds for the training and validation data decay as expected to a point of stability with a minimal gap between the two final loss values as shown in Fig. 4.
Figure 4: Training and validation loss curves
The mean squared error (MSE), mean absolute error (MAE) and $R^{2}$ score performance metrics for training, validation, and test data demonstrate the prediction performance of the model. These metrics are given by Eqs. 2-4.
${MAE}=\frac{1}{N}{\sum_{i=1}^{N}}\left|y_{i}-\hat{y}\right|$ (2)
${MSE}={\frac{1}{N}{\sum_{i=1}^{N}}(y_{i}-\hat{y})^{2}}$ (3)
${R^{2}}={1-\frac{{\sum_{i=1}^{N}}(y_{i}-\hat{y})^{2}}{{\sum_{i=1}^{N}}(y_{i}-\bar{y})^{2}}}$ (4)
where $N$ is the number of data points, $\hat{y}$ is the predicted value of $y$, $y_{i}$ is the values of $y$ variable in the sample, and $\bar{y}$ is the mean value of $y$. The mean squared error (MSE), mean absolute error (MAE) and $R^{2}$ score for training, validation, and test data of the target features are tabulated in Table IV. We note that features $F_{x}$ and $F_{y}$, appearing on the table, represent the contact forces in the x- and y-directions, respectively. In a similar vein, features $D_{x}$ and $D_{y}$ indicate the tip displacements, respectively, in the x- and y- directions. These numbers suggest that the model is suitably generalized to unseen data and our model has a balance between bias and variance, avoiding overfitting and underfitting. Therefore, based on the performance metrics and numerical results presented here, the MLP is capable of precise predicting of the target features for various internal structures.
TABLE IV: MSE, MAE, and $R^{2}$ score performance metrics
Error Type Feature MSE MAE $R^{2}$ score
Training $F_{x}$ 0.00084 0.02245 0.99
Validation $F_{x}$ 0.00068 0.01954 0.99
Test $F_{x}$ 0.00104 0.02537 0.99
Training $F_{y}$ 0.00351 0.04156 0.94
Validation $F_{y}$ 0.00217 0.03760 0.94
Test $F_{y}$ 0.00143 0.02804 0.98
Training $D_{x}$ 0.00338 0.04282 0.94
Validation $D_{x}$ 0.00177 0.03141 0.95
Test $D_{x}$ 0.00254 0.04111 0.95
Training $D_{y}$ 0.00241 0.03562 0.96
Validation $D_{y}$ 0.00184 0.02823 0.96
Test $D_{y}$ 0.00198 0.03707 0.97
IV Multi-objective Design Optimization
We employ a multi-objective optimization algorithm to find the optimal internal structure for the Fin-Ray finger. We use NSGA-II multi-objective genetic algorithm, which generates offspring by using a particular kind of crossover and mutation method. The selection process for the next generations is based on comparisons of the non-dominated sorting and crowding distances[21]. Three design parameters are considered in the dimensional optimization, including the thickness of the front and supporting beams, the thickness of the crossing beams, and the equal spacing between the crossbeams, which are considered input features for the MLP in the previous section. These design parameters are constrained in the range that the MLP is trained according to Table I. The MLP model developed in the previous section is used here for accurate prediction of our objectives.
Two objectives of optimization are readily determined from the four target features of the MLP using Eqs. 5 and 6:
${F}={\sqrt{F_{x}^{2}+{F_{y}^{2}}}}$ (5)
${D}={\sqrt{D_{x}^{2}+{D_{y}^{2}}}}$ (6)
where $F$ and $D$ are the predicted magnitude of the maximum contact force and magnitude of the maximum tip displacement, $F_{x}$ and $D_{x}$ are the predicted maximum force and maximum tip displacement in the x-direction, and $F_{y}$ and $D_{y}$ are the predicted maximum force and maximum tip displacement in the y-direction, respectively. As aforementioned, these magnitudes are considered as our two objectives. The maximum tip displacement towards the grasped object is regarded a scale for how suitable a design is at handling delicate objects and highlights the Fin-Ray finger’s adaptability to the cylindrical objects. A higher value for the contact force resembles the tendency of the finger to possess a higher load capacity. Achieving a large grasping force demands enhanced rigidity, which is in contrast with the first objective and it is against the softness, necessary for delicate grasping.
After the determination of the objectives and design parameters, NSGA-II is utilized to find the optimal design of the Fin-Ray finger. The Pymoo library allows to implement the NSGA-II algorithm by defining our problem with the specific parameters mentioned in Table V[22]. The set of optimal solutions, which is also known as the Pareto front, is presented in Fig. 5. Pareto front narrows down the solutions to a set in which we can make choice in a trade-off of one objective over the others. Depending on application, one of these solutions is selected to achieve desirable design.
TABLE V: NSGA-II algorithm parameters
         Parameter Name          Value
         Number of generations          100
         Population size          500
         Crossover rate          0.9
         Mutation rate          0.1
Figure 5: The set of all Pareto efficient solutions
We also conduct a series of numerical experiments, in which we feed the MLP random numbers and compare the predicted target features with the Pareto front. However, the Pareto front is not dominated by the experiment data or dataset as illustrated in Fig. 6.
Figure 6: Experiment data, dataset, and Pareto front
V Conclusion
In this research, we have presented an comprehensive approach for designing the Fin-Ray finger to enhance the grasping performance. A finite element framework for predicting the forces during the gripper’s interaction with cylindrical objects has been proposed. Using finite element simulations, we have developed a dataset consisting of distinct Fin-Ray internal structures. A model has been forged through training the neural network. The chosen MLP hyperparameters have been selected via a grid search over the specified range of values. This MLP model has been used to find the optimal structures with the NSGA-II algorithm. The set of all Pareto efficient solutions has been found and validated with the experiment data and the dataset itself. Hence, it is found that the present methodology achieved the desired results for designing the optimal Fin-ray finger, assisting to make a choice in the trade-off of one objective over the other.